# Fair Overlap Number of Balls (Fair-ONB): A Data-Morphology-based Undersampling Method for Bias Reduction

### v1

@@ 6 Concluding remarks
This paper shows the usefulness of morphology and neighbourhood-based sampling approaches towards classifier fairness, a niche that has not been thoroughly explored. Fair-ONB, which employs guided undersampling in areas that are close to the decision boundaries, where models usually have problems classifying the different protected groups. The Fair-ONB method harnesses the morphology of groups to select samples for elimination, and unlike other random sampling techniques, its logic is more justifiable due to its empirical approach.
The goodness of the Fair-ONB method is demonstrated in two steps. The first study checks its behaviour on 4 datasets, where it was able to produce better fairness results in all cases, accompanied by better classification performance results in half of them. While no clear pattern was obtained regarding the threshold variables, it was shown that they provide a proper characterisation of the problematic areas in datasets and, thus, that using subgroup ball coverage for undersampling instance selection is a very useful strategy.
The second study checks that the Fair-ONB method is able to better bolster fairness compared to FAWOS, a similar state-of-the-art neighbourhood-based sampling method. Not only does the Fair-ONB method have lower computational complexity, but its identification of the problematic areas, which takes into account more than just a few neighbours, allows it to reach better results.
Finally, regarding future work lines, there are multiple possible courses of action.
  • Employing the group characterisation of the Fair-ONB method on datasets that do not contain protected variables (in which case, groups would be given by just the classes). This way, noise-reducing preprocessing could be applied to improve classification performance in situations with poor data quality.
  • Combining morphology-based strategies with other preprocessing techniques, so that higher fairness levels can be attained. Using a bigger scope could lead to finding synergies and that could lead to more in-depth preprocessing.
@@ Acknowledgments
This work has received funding from the Spanish Ministry of Science and Technology under project PID2020-119478GB-I00, including European Regional Development Funds, and the “Cátedra Tecnalia en Inteligencia Artificial” programme. It is also partially supported by the I+D+i project granted by C-ING-250-UGR23 co-funded by “Consejería de Universidad, Investigación e Innovación” and the European Union related to FEDER Andalucía Program 2021-27.

### v2

@@ 6 Concluding remarks
This paper presents Fair-ONB, a new sampling method that employs guided undersampling in areas that are close to the decision boundaries, where models usually have problems classifying the different protected groups. The Fair-ONB method harnesses the morphology of groups to select samples for elimination, and unlike other random sampling techniques, its logic is more justifiable due to its empirical approach.
As shown in the experimental study, the Fair-ONB method improves model Fairness while maintaining or improving classification performance, reflecting the usefulness of using subgroup ball coverage for undersampling instance selection. Moreover, the Fair-ONB method also achieves better bias reduction results than FAWOS, its neighbourhood-based competitor from the state of the art.
Therefore, this paper has shown the usefulness of morphology and neighbourhood-based sampling approaches towards classifier Fairness, a niche that has not been thoroughly explored. In particular, the use of threshold variables and group coverage strategies, which give information on multiple morphological properties of the groups in a dataset, have been useful for the proper characterisation of problematic overlap areas. Thus, the hypothesis that using data morphology for undersampling instance selection can be a very useful strategy is confirmed.
Regarding future work lines, combining morphology-based strategies with other preprocessing techniques could lead to attaining higher fairness levels. Using a bigger scope and finding synergies could allow for the implementation of more in-depth preprocessing techniques.
@@ Acknowledgments
This work has received funding from the Spanish Ministry of Science and Technology under project PID2020-119478GB-I00, including European Regional Development Funds. It is also partially supported by the I+D+i project granted by C-ING-250-UGR23 co-funded by “Consejería de Universidad, Investigación e Innovación” and the European Union related to FEDER Andalucía Program 2021-27.